OpenStudy (helpkay):

please help @rishavraj @bighungy270

OpenStudy (helpkay):

do you do 2(20+ 4), find or 2(20)+ 4),

OpenStudy (helpkay):

@bighungry270 @rishavraj

rishavraj (rishavraj):

so it simply means 20 = 2(x + 4)

OpenStudy (helpkay):

i know that but you have to plug 20 in for both of the X's so wich way do you plug it in

rishavraj (rishavraj):

wht do u mean ....u just need to solve the equation i mentioned above ..... u will get x

OpenStudy (helpkay):

I can only give you the text version of what I am dingo but I know that you have to solve the problem is solving it

OpenStudy (helpkay):

If f of x equals three x plus one, find f of negative 2. The screen shows the function described on line one with the problem statement next to it. You are being asked to find the value of the function when x is equal to negative 2. That is, the value of x is explicitly given to you, and you must substitute each instance of x in f of x with its equivalent, negative 2. Once you finish that, all that's left to do is simplify. Then, the function from line one is copied down to the middle of line two. Each time you see x in the function rule, substitute negative 2. Then, the function from line two is copied down to line three. The x's from f of x and 3 x are briefly highlighted. Then, the x's are morphed into negative 2s. So, the function on line three becomes f of negative 2 equals 3 times negative 2 plus 1. Simplify the right side of the function, keeping in mind the order of operations. Then, the function from line three is copied down to line four. The expression 3 times negative 2 is morphed into negative 6. So the function becomes f of negative 2 equals negative 6 plus 1. Then, the function from line four is copied down to line five. The expression negative 6 plus 1 is morphed into negative 5. So, the final function is f of negative 2 equals negative 5.

OpenStudy (anonymous):

hold on

rishavraj (rishavraj):

tell me onething.... if a = 3x and a = 18 it means 3x = 18 agreed??

OpenStudy (helpkay):

yes

OpenStudy (helpkay):

that was an example I sent you here is the screen shot of the example

rishavraj (rishavraj):

in the recent pic u r given the value of x ...bt te question u posted is reverse u need to find the 'x'

rishavraj (rishavraj):

so like f(x) = 2(x + 4) and f(x) = 20 means 2(x + 4) = 20

OpenStudy (helpkay):

see you have to plug in -2 for both X that what I need help with in this question how do i plug the number in for x in 2(x + 4),

OpenStudy (helpkay):

sorry didn't see your post

rishavraj (rishavraj):

u r moving the wrong direction.... u need to find x which satisfies 2(x+ 4) = 20

OpenStudy (helpkay):

that is not what they teach me they say you plug in 20 in the place of x for f(x) = 2(x + 4), wich would be f20 = 2(x + 4), then you have to plug it in for the X in 2(x + 4), wich I think would be 2(20) + 4), but i'm not 100% sure I will send you an example

OpenStudy (helpkay):

how i solved this is you do h(x) = −5x − 10, find h(−2) so you plug in -2 for both Xes h-2=-5(-2)-10 h-2=-5 X -2-10 h-2=10-10 h-2=0 @rishavraj

rishavraj (rishavraj):

its h(-2) = -5(-2) - 10 h(-2) = 0 the thing is tht here u r given the value of x ..and asked the value of the function at tht point ...bt in ur question (the posted one) u are having the value of funtion at some "x" and askd to find out tht 'x'

OpenStudy (helpkay):

oh i didn't see that so there not asking you to solve there asking you to find x that why I was so confused

rishavraj (rishavraj):

yup :)) now u got tht

OpenStudy (helpkay):

so it would be 8

OpenStudy (helpkay):

@rishavraj is it 8

rishavraj (rishavraj):

why 8 ??guessing??? see 20 = 2(x + 4) wht a(x + y)

rishavraj (rishavraj):

lol I meant *whats a(x + y)

OpenStudy (helpkay):

? wht a(x + y) no wait it would be 6 because you have to multiply both numbers

rishavraj (rishavraj):

yup 6 is correct :))

OpenStudy (helpkay):

thanks for all you help I was so confused you deserve 5 medals for all that help @rishavraj

rishavraj (rishavraj):

lol u welcome :))) @helpkay